OpenStudy (anonymous):

what is the point-slope form of the equation of a line passes through the point 6,-8 and has a slope of -2 A.y+8=2(x-6) B.y+8=-2(x-6) C.y-6=-2(x-8) D.y-8=-2(x+6)

OpenStudy (freckles):

big hint the point slope form of a line that passes through (a,b) with slope m is y-b=m(x-a)

OpenStudy (anonymous):

oh i know C i think @freckles ?

OpenStudy (anonymous):

m is slope & it states that the slope is -2 freckles said: y-b=m(x-a) that means its either B, C, or D

OpenStudy (freckles):

@mamas (a,b)=(6,-8) in this case I think you mixed b and a around

OpenStudy (anonymous):

(a,b) is (6,-8) so a =6 and b = -8 so you have: y-b = -2(x-a) y+8=-2(x-6)

OpenStudy (anonymous):

B, y + 8 = -2(x-6)

OpenStudy (anonymous):

thats the answer @mathrulezz ? @freckles ? B

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thanks @mathrulezz

OpenStudy (anonymous):

np